METCRAX: Logbook Entries

METCRAX: Soil Messages: 9 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
188 Thu 09-Aug-2007Qsoil adjustmentsAllOncley
187 Wed 08-Aug-2007gravimetric summaryAllOncley
186 Wed 23-May-2007Change of Qsoil calibration to mineral soilAllHorst
172 Fri 22-Dec-2006Soil are quality controlledAllaster
161 Wed 01-Nov-2006flr TP01FloorOncley
155 Mon 30-Oct-2006flr TP01 never hooked upFloorOncley
144 Fri 27-Oct-2006soil reinstallationAllSemmer
108 Sat 21-Oct-2006gravimetric measurementsOncley
82 Sun 15-Oct-2006dirtBaseMartin

188: Soil, Site All, Thu 09-Aug-2007 15:51:48 MDT, Qsoil adjustments
Based on the manual measurements (gravimetric+TRIME) in the previous entry,
I've adjusted biases on the Qsoil measurements, which match to within typically
1%.  Note that we don't have enough variability in the manual measurements to
determine if the slope should be corrected:

flr	no correction
el	+5%
wl	+4%
eu	+5%
wu	+5%
rim	no correction
sw	-5%

187: Soil, Site All, Wed 08-Aug-2007 16:24:39 MDT, gravimetric summary
Here is my reduction of the manual soil measurements:

Site	Date	Time	Rho	Grav	Rho*Grav	Trime
			g/cm3	%	%vol		%vol
flr	10/21	1315	1.128	5.5	6.2		-
eu	10/21	1332	1.496	3.5	5.2		-
wu	10/22	1502	1.111	3.0	3.3		5.7
wl	10/22	1517	1.317	1.3	1.7		2.2
el	10/22	1540	1.167	1.0	1.2		3.7
eu	10/22	1555	1.927	3.8	7.3		8.2
rim	10/24	1600	1.553	4.6	7.1		4.3
sw	10/24	1700	1.481	10.5	14.9		13.4
wu	10/26	1154					5.4
wl	10/26	1210					2.9
flr	10/26	1224					6.6
el	10/26	1237					3.0
eu	10/26	1253					6.3
rim	10/30	1500					3.4

186: Soil, Site All, Wed 23-May-2007 12:02:35 MDT, Change of Qsoil calibration to mineral soil
Alan suggests changing the Qsoil calibration from 
Qsoil = 0.000695*Vqsoil(mV) - .29
to that for mineral soils, which is
Qsoil = 0.00119 * Vqsoil - .401.

Thus the archived data should be changed by = 1.7122 * Q.old + 0.0955

[Actually, this is wrong -- this is for Qsoil as a fraction, but we save Qsoil
in %.  Thus, the correction should be: = 1.7122*Q.old + 9.55....Steve O.]

[Furthermore, this appears to be a misread of the Ech2o manual.  I believe that
the top equation is correct for the EC-20 probes that we used (in any soil).
Nevertheless, I've adjusted biases based on manual measurements -- see later
comments...Steve O.]

172: Soil, Site All, Fri 22-Dec-2006 14:11:08 MST, Soil are quality controlled
Soils have been quality controlled

1. Gsoil, note: Gsoil.flr had some bad data removed at the beginning of project, Gsoil.rim had some bad data glitches around Oct. 25, everything else is ok.
2. Tsoil, note: Tsoil.flr had some bad data removed at the beginnin gof project,
everything else ok except some random outliers removed.
3. Qsoil, note: need to apply new mineral soil calibrations to replace factory
calibrations; most of the data is ok except there are glitches in Qsoil.rim 
after Oct.24, removed a fair amount of data.
4.Cvsoil.flr dead entire project ok
  Cvsoil.el ok
  Cvsoil.rim ok
  Cvsoil.wl ok
  Cvsoil.wu ok
  Cvsoil.sw ok

161: Soil, Site Floor, Wed 01-Nov-2006 16:31:39 MST, flr TP01
At about 0900 yesterday, I removed the TP01 cabling from sw.
At about 1030 yesterday, I reinstalled it in flr's logger and we started getting
data for the first time.  The values appear to be consistent, so I think this
will give values that can be applied back to the 6 Oct rain event.

155: Soil, Site Floor, Mon 30-Oct-2006 16:20:31 MST, flr TP01 never hooked up
We have just discovered that the wiring to the Amp connector inside the logger
box for the TP01 is completely missing, thus we have never acquired data from
this sensor for METCRAX.  My guess is that the wiring might have been removed
from this box to fix another box during either TREX or CuPIDO and never 
replaced.  We plan to put another logger box at this site tomorrow and thus
get 2 days of good data before flr's teardown.

144: Soil, Site All, Fri 27-Oct-2006 11:47:45 MST, soil reinstallation
Tom asked that this be put into the records:
During the installation phase there was the need to move some of the 
sawhorses. This required the reinstalltion of soem soil sensors.
The following was done.
  On Sept 29 the floor soil samples were install
  On Sept 28th, the soil sensors at West Upper were moved and reinstalled.
  No other sensors were removed and installed due to the change in the sawhorse
108: Soil, Site , Sat 21-Oct-2006 15:32:53 MST, gravimetric measurements
I took the first gravimetric measurements today.

Since we couldn't find the brass rings, I simply used the whole corer, which
samples deeper.  With a tape measure, I tried to estimate the depth of the
core (I didn't quite get either core full).  The corer without rings is about
2 1/4" diameter.

Also note that, since the TRIME battery was dead, no measurements could be
made.  It is now charging back in the base.

The weight of the cores exceeds the scale range (200g), so they have been
arbitrarily split into 2 trays, hence the sum of 2 numbers below.

flr: core extracted ~1315 10/21, core depth ~3 7/8"
     total weight: 130.8+165.9g = 296.7
     dry: 124.5+156.7g = 281.2 (5.5%)

eu:  core extracted 1332 10/21, depth ~2 3/8" [hit a rock, so couldn't go 
     total weight: 122.9+116.8g = 239.7
     dry: 118.6+112.9g = 231.5 (3.5%)

Measurements taken 10/22:
[Found brass rings, so these are 0-6cm samples.  Note that the core often wasn't
full to the top when taken apart, but I'm guessing that the mass is about right
and that the soil compacted upon being cored.  Nevertheless, there usually was
 spillage, so bulk densities would be slightly underestimated.]
[Also had TRIME charged, so took readings coincident with the cores.]
wu:  time: 1502, TRIME: 5.7%
	total: 153.3g
        dry: 148.9g 

wl:  time: 1517, TRIME: 2.2%
	total: 114.0+64.9g = 178.9g
        dry: 112.5+64.1g = 176.6g

el:  time: 1540, TRIME: 3.7%
	total: 158.1g
        dry: 156.5g

eu:  time: 1555, TRIME: 8.2%
	total: 127.5+140.7g
        dry 122.7+135.7g = 258.4g

Measurements 10/24:
[hauled the TRIME and corer up from the crater to do these,
 note that we had a bit of rain today before these measurments were made]
rim:  time: ~1600, TRIME: 4.3%
      total: 112.4+105.4g
      dry: 107.2+101.0g
sw:   time: ~1700, TRIME: 13.4%
      total: 88.9+121.2g
      dry: 81.0+109.1g

Measurements 10/26:
wu:   time 1154, TRIME: 5.4%
wl:   time 1210, TRIME: 2.9%
flr:  time 1224, TRIME: 6.6%
el:   time 1237, TRIME: 3.0%
eu:   time 1253, TRIME: 6.3%

Measurement 10/30:

rim:  time ~1500, TRIME 3.4%

82: Soil, Site Base, Sun 15-Oct-2006 17:10:18 MST, dirt
I have loaded Greg's Cupido dirt into my car
and will deliver it to him when I get back
to Boulder.